hey maybe someone could help me with these three algebra problems, i have been sitting on them almost all day, and my homework is do by monday (ill get a D if i dont give my homework up this time:( )so its sorta important to get them done, but i dont understand them :/ so enough of the small talk ill just tell you which algebra problems they are;
1) find an equation of the line point slope form that passes through (4,3) and is parallel to the line 3x-y=7
2) what is the equation in standard form of the line through (4,1) an (-2,3)?
3) find the equation of the line with x-intercept 5 and y-intercept 8
it would be great if someone could help me with these! :)

Answer by Alan3354(69443)

1) find an equation of the line point slope form that passes through (4,3) and is parallel to the line 3x-y=7
Find the slope of the line by putting 3x-y=7 in slope-intercept form.
3x-y=7
y = 3x - 7 so the slope, m, is 3.
A parallel line will have the same slope.
Use y-y1 = m*(x-x1) where (x1,y1) is the point (4,3)
y-3 = 3*(x-4)
y-3 = 3x-12
y = 3x-9
-------------
2) what is the equation in standard form of the line through (4,1) an (-2,3)?
The hard way to do this:
1st, find the slope, m
m = (diff in y)/(diff in x)
m = (1-3)/(4-(-2)) = -2/6
m = -1/3
Then use either point, same as in #1.
y-1 = (-1/3)*(x-4)
y-1 = -x/3 + 4/3
y = (-1/3)x + 7/3 slope-intercept form
-----------------
x + 3y = 7 standard form
-----------------------
3) find the equation of the line with x-intercept 5 and y-intercept 8
That means it goes thru the points (5,0) and (0,8)
Same as problem #2
Try it, email me for help or to check your work @ gsihoutx@aol.com